OpenStudy (anonymous):

y=2x is a tangent to the curve y=x^3+ax+b at x=1. Find a and b.

OpenStudy (bahrom7893):

Tangent line is the first derivative, so: y' = 3x^2 + a; x = 1; y' = 3*1^2 + a

OpenStudy (bahrom7893):

3 + a = y'

OpenStudy (bahrom7893):

now slope of ur tangent line is: y= 2x; M = 2=> 3 + a = 2 => a = -1

OpenStudy (bahrom7893):

but what the heck is b? lol im lost lemme see, kinda embarassing

OpenStudy (anonymous):

wouldn't i just fill in a & fill in 1 for x to find b?

OpenStudy (bahrom7893):

oh yeah, i think you do this: you need y too when x=1, so y = 2x = 2; x = 1; a = -1 2 = 1^3 + (-1)*1 + b 2 = 1 -1 +b => b = 2

OpenStudy (bahrom7893):

can u check the answer?

OpenStudy (anonymous):

yes that's right thank you so much!